The standard labor cost in the production of a pair of Beca Brand athletic shoes is 0.40 hours at $14.00 per hour. During the month of June, 24,630 pairs were produced. Actual labor costs were $96,240 for 8,020 hours. Compute the labor rate and labor efficiency variances for the month of June. (Round actual wage rate to 2 decimal places, e.g. 15.25 and final answers to 0 decimal places, e.g. 125. Enter all variances as a positive number.)

Labor Rate Variance $

Neither Unfavorable nor FavorableFavorableUnfavorable

Labor Efficiency Variance $

UnfavorableFavorableNeither Unfavorable nor Favorable
Answer #1

Labor rate variance = (SR-AR) *AH

= (14 - 96,240/8020)*8020

= 16,040 Favourable

.

Labor efficiency variance = (SH-AH)*SR

= (24,630*0.40 - 8020)*14

= 25,648 Favourable